OpenStudy (anonymous):

Two right triangles with the following dimensions are graphed with their hypotenuses on the same line: triangle X: Horizontal leg measures 2 units vertical leg measures 5 units triangle Y: Horizontal leg measures 7 units vertical leg measures 17.5 units Which other triangles could also have their hypotenuse on the same line? Choose all answers that are correct. A. horizontal leg measures 8 units, vertical leg measures 10 units B. horizontal leg measures 4 units, vertical leg measures 10 units C. horizontal leg

myininaya (myininaya):

Think about slope.

myininaya (myininaya):

vertical/horizontal

OpenStudy (anonymous):

I have . I even bloody graphed the damn points and nothing adds up @!!@!

myininaya (myininaya):

5/2=17.5/7=2.5

myininaya (myininaya):

see which of the others give the same ratio